# Detecting Masquerade Attacks in Controller Area Networks Using Graph Machine Learning

### v1

@@ III-A Prior Work Closely Related to the Present Study
Islam et al. [43] proposed a four-stage IDS utilizing the chi-squared method to identify both strong and weak cyberattacks on the CAN bus. Their approach, which is among the first graph-based defense systems for CAN, showed misclassification rates of 5.26% for DoS attacks, 10% for fuzzy attacks, 4.76% for replay attacks, and zero misclassification for spoofing attacks. The dataset used for evaluation was obtained from the Hacking and Countermeasure Research Lab [44]. This study emphasized the need for robust security mechanisms in modern vehicles and demonstrated superior accuracy compared to existing ID sequence-based methods.
@@
Jedh et al. [21] proposed a message injection attack detection solution. Their approach leveraged MSGs and used graph-based analytics and anomaly detection to detect malicious message injections with high accuracy and low detection time. They validated their approach using a dataset collected from a moving vehicle under attack conditions. This study highlighted the importance of addressing the security of in-vehicle communication networks without relying on proprietary ECU information.
Islam et al. [45] introduced a novel graph-based Gaussian Naive Bayes (GGNB) algorithm for CAN IDS. They leveraged graph-based features from the nodes and edges. The GGNB method applied to the real raw CAN dataset, achieved high detection accuracy across various attack types, including DoS, fuzzy, spoofing, and replay attacks. The study also utilized the OpelAstra dataset [46] and reported significant improvements in training and testing times compared to SVM classifiers. This research showed the efficiency and effectiveness of the GGNB-based methodology in a resource-constrained environment.
Sreelekshmi et al. [47] proposed a graph-based IDS for CANs achieving notable accuracy in detecting various types of attacks. Their method leverages bidirected graphs and parameters like degree variance to identify anomalies. The system was tested on the Car Hacking: Attack & Defense Challenge 2020 dataset [48] demonstrating superior performance with an accuracy of 98.38%.
@@
Refat et al. [49] presented a lightweight IDS that translates CAN traffic into temporal graphs and applies neighborhood-based graph similarity techniques to detect message injection attacks. The system was evaluated using real vehicle data [50], achieving a detection accuracy of 96.01% for spoofing, fuzzy, and DoS attacks. This work contributed to vehicle security by providing a computationally efficient method without requiring changes to the CAN protocol.
@@
Zhang et al. [23] proposed a CAN bus anomaly detection system using graph neural networks (GNNs) to detect message injection, suspension, and falsification attacks. Their approach involved creating directed attributed graphs from CAN message streams and training a two-stage classifier cascade. The evaluation on a Ford Transit 500 dataset [51] demonstrated the system’s efficiency in real-time detection and its ability to handle new anomalies through federated learning training.
Park et al. [52] introduced the G-IDCS, a graph-based IDS integrating a threshold-based IDS and a ML-based attack classifier. This system reduced the number of required CAN messages for detection and improved detection accuracy by over 9% compared to existing methods. They used the Car Hacking: Attack & Defense Challenge 2020 dataset [48] for evaluation. Their study emphasized the system’s robustness to changes in attack types and its potential use in digital forensic investigations.
@@
Xiao et al. [14] proposed the CAN-GAT model based on graph attention networks for in-vehicle networks. By transforming CAN bus messages into graph structures, the model captured the correlation between traffic bytes, improving detection accuracy and efficiency. The model was evaluated using datasets from the Hacking and Countermeasure Research Lab [44], demonstrating superior performance among compared GNNs.
@@
Meng et al. [53] developed GB-IDS, an IDS leveraging a novel graph structure and a variational autoencoder for training classifiers without negative samples. Their system, evaluated on the OTIDS dataset [54], achieved high detection success rates for DoS, fuzzing, and impersonation attacks. This study addressed the limitations of traditional IDS, i.e., [49] by avoiding the need for protocol parsing and large training datasets.
In this paper, we propose a unique framework for detecting masquerade attacks in CANs, leveraging graph ML and time series analysis. By representing CAN frames as MSGs and annotating their nodes with time series features, we fill a gap in the current research. Our approach builds upon previous graph-based methods, such as those introduced by Islam et al.[43], Jedh et al.[21], and demonstrates competitive performance against recent studies using the ROAD dataset [31] [55]. Unlike methods focusing solely on time series data, our framework captures both structural and temporal aspects of CAN frames, enhancing detection capabilities. The flexibility of our framework, adjustable via the window length and offset in a sliding window environment, allows for fine-tuning performance in constrained vehicular systems filling the gaps in [23]. We test our framework on the most realistic masquerade attacks from the ROAD Dataset [15], demonstrating its potential for enhancing automotive cybersecurity.
@@ III-B Other Prior Work Related to the Present Study
Zhou et al. [56] developed BTMonitor, a bit-time-based IDS for CANs, achieving a sender identification accuracy of 99.76% on a real vehicle. Their approach leverages the small discrepancies in bit times to fingerprint sender ECUs and can detect new types of masquerade attacks that other IDS fail to identify. They evaluated the system using a three-node CAN bus prototype and a 2012 Buick Regal production vehicle dataset.
Ying et al. [57] proposed a formal analysis of clock skew-based IDS and introduced the cloaking attack, a new type of masquerade attack that manipulates message inter-transmission times to avoid detection. Their study, validated using hardware testbeds and the UW EcoCAR dataset [58], predicted the attack success probability with an average prediction error of within 3.0%.
Hanselmann et al. [59] proposed CANet, a neural network architecture for detecting intrusions on the CAN bus. CANet, the first deep learning-based IDS capable of handling the high dimensionality of CAN bus data, was evaluated on real and synthetic CAN data. The method outperformed previous ML-based approaches with a high true negative rate, typically over 0.99, demonstrating robustness in detecting a large number of synthetic masquerade attacks [59].
Moriano et al. [31] focused on detecting masquerade attacks on the CAN bus by analyzing time series extracted from raw CAN frames. Using hierarchical clustering, this study demonstrated that changes in cluster assignments could indicate anomalous behavior. The proposed forensic tool was tested on the ROAD dataset [15], showing significant differences in time series clustering similarity on benign and attack conditions.
Shahriar et al. [55] developed CANShield, a deep learning-based signal-level intrusion detection framework for the CAN bus. CANShield handles high-dimensional CAN data streams, analyzing time-series data from different temporal scales using multiple deep autoencoder networks. Evaluation on the SynCAN [59] and ROAD [15] datasets demonstrated the framework’s robustness against various advanced attacks, improving the overall AUC-ROC by 6.40% compared to conventional methods.
Sharmin et al. [8] contributed to CAN IDS benchmarking efforts by comparing six CAN IDS algorithms using the ROAD dataset [15]. The study included algorithms based on ID sequences, entropy, Hamming distance, frequency and isolation forest. Their results showed that entropy- and frequency-based algorithms performed better emphasizing the need for fine-tuning existing methods to detect sophisticated attacks like targeted ID and masquerade attacks.
Shahriar et al. [3] proposed CANtropy, a feature engineering-based lightweight CAN IDS. CANtropy explores a comprehensive set of features from both temporal and statistical domains utilizing PCA for anomaly detection. Evaluation on the SynCAN dataset [59] showed CANtropy’s effectiveness, with an average AUC-ROC score of 0.992, outperforming existing DL-based baselines.
Moriano et al. [60] conducted a benchmark study on four non-deep learning unsupervised online IDS for masquerade attacks in CANs. They controlled streaming data conditions in a sliding window setting and used realistic attacks from the ROAD dataset [15]. They noticed that there is no one-size-fits-all solution for detecting masquerade attacks in CAN in an online setting, i.e., there is no single algorithm that is optimal for detecting each attack type. Notably, a method that detects changes in the hierarchical structure of clusters of time series outperformed others in detecting various attack types at the expense of computational overhead.
@@ IV Proposed data-driven Methods
This section introduces our proposed approach for detecting masquerade attacks in the CAN bus using graph ML. We begin by defining key terms and then go into the details of our processing pipeline illustrated in Fig. 3. We first discuss the partitioning of CAN data for building MSGs (Section IV-A), followed by the extraction and analysis of time series data (Section IV-B). Next, we detail the process to annotate the graphs (Section IV-C) and the generation of graph embeddings (Section IV-D). We then describe our supervised learning process (Section IV-E). Finally, we outline our evaluation setup and metrics (Section IV-F).
@@ IV-A CAN Data Partitioning and MSG Building
We use sliding windows of fixed size to partition and process the stream of CAN data [61]. The raw CAN data is first transformed into a structured dataframe format where each row represents a message with its timestamp, process ID, and data payload (see Fig. 4). This data is then partitioned into discrete time slices defined by specific window sizes and offsets, enabling a continuous analysis of CAN data.
@@
  • Window size ($\omega$): The length of each window determining the extent of data included in each graph representation.
  • Offset ($\delta$): The sliding step between consecutive windows, where $\delta$ represents the number of samples separating each window, ensuring no loss of data and the capture of communication patterns that may span multiple windows.
A MSG is constructed within each defined sliding window to model CAN activity following the procedure introduced by [21] and [43]. Graphs are empirically recognized as effective tools to model relationships among relational data that are too complex for tabular representation or other simpler data structures [62]. The process of building MSGs is crucial as it sets the stage for subsequent analyses, including node annotation and embedding generation. By modeling CAN data using MSG on sliding windows, we enhance our ability to detect and respond to potential masquerade attacks by finetuning the portion of analyzed data. We now detail each of the components of the MSG:
@@
  • Edges: Directed edges are established based on the sequential relationships of messages. That is, an edge is formed when one ID follows another, indicating the flow of communication [43]. This method effectively maps sequences of CAN messages enabling the detection of irregular patterns that may signify an attack. Edge weights are assigned to edges of MSGs patterns to quantify the frequency of observed communication sequences within a sliding window. Specifically, the weight of an edge from node $v_{i}$ to node $v_{j}$ is computed as the number of times the CAN message with ID $v_{j}$ directly follows the CAN message with ID $v_{i}$ within the same sliding window. Weights help to identify anomalous recurrent patterns that could indicate security threats.
Figure 4: Sliding windows concept. The left side shows raw CAN bus messages partitioned by a window of length $\omega$ and offset $\delta$. The right side depicts a resulting MSG subgraph, highlighting the top five nodes with the highest degrees of connectivity within the first sliding window. Edges are annotated with weights.
@@
0:  Raw CAN data, $\omega$, $\delta$
@@
2:  Initialize $current\_time=start\_time$ of CAN data
3:  while $current\_time+\omega\leq end\_time$ of CAN data do
4:     Define $start\_time=current\_time$ and $end\_time=current\_time+\omega$
5:     Extract messages in $[start\_time,end\_time]$
@@
7:     for each message $m$ in $[start\_time,end\_time]$ do
8:        Extract $id$ from $m$
9:        if $id$ not in $G_{sub}$ then
10:           Add $id$ as a node in $G_{sub}$
@@
12:        Determine next $id$ in sequence
13:        if edge from current $id$ to next $id$ exists then
@@
20:     Increment $current\_time$ by $\delta$
@@
To ensure the correctness of Algorithm 1, we establish a loop invariant for the main while loop. A loop invariant is a property that holds true before and after each iteration of the loop, helping to trace the algorithm’s functionality and correctness. We demonstrate that the loop invariant holds through three key steps:
@@
    Loop Invariant: At the start of each iteration of the while loop, current_time is a valid start time for a new window, and all CAN messages up to current_time have been processed into the main graph $G$.
    • Initialization: The invariant holds when current_time is set to the start time of the CAN data. No messages have been processed yet, and $G$ is empty. This ensures that the loop invariant is true before the loop starts.
    • Maintenance: During each iteration, a window is defined from current_time to current_time + $\omega$. All messages in this window are extracted and processed into a subgraph $G_{\text{sub}}$, which is then merged into the main graph $G$. current_time is incremented by $\delta$, updating it to a new valid start time for the next window, ensuring all messages up to the new current_time have been processed. This demonstrates that the property holds after each loop iteration.
    • Termination: The loop terminates when current_time + $\omega$ exceeds the end_time of the CAN data. At this point, there are no more complete windows of size $\omega$ that can start at or after current_time. The invariant holds true because current_time has moved past the end of the data, confirming that all messages in the valid windows have been processed. The loop invariant confirms that the algorithm functions as intended and supports its correctness.
    By maintaining the loop invariant throughout the execution, we demonstrate the correctness and robustness of the MSG creation algorithm.
@@
    Time Complexity: The time complexity of our MSG creation algorithm depends on $\omega$ and $\delta$. In general, it can be expressed as $O\left(\frac{n}{\delta}\cdot\omega\right)$, where $n$ is the total number of messages in the CAN data. This is calculated as follows: The outer loop in Algorithm 1 iterates approximately $\frac{n}{\delta}$ times, processing the CAN data in steps of $\delta$. Within each iteration, a window of size $\omega$ is processed, involving operations such as extracting messages and constructing subgraphs, which take $O(\omega)$ time. Therefore, the overall time complexity is the product of the number of iterations and the operations per iteration, leading to $O\left(\frac{n}{\delta}\cdot\omega\right)$.
    In our experiments, we test various combinations of $\omega$ and $\delta$, ranging from (2, 1) to (15, 14) for time-based windows and from (50, 50) to (400, 400) for sample-based windows. For each attack, we run the algorithm with these combinations to analyze the impact of $\omega$ and $\delta$ on performance and detection accuracy. For example, with a combination of $\omega=4$ and $\delta=4$, the time complexity simplifies to $O(n)$, as each message is processed exactly once. This reflects the trade-off between $\omega$ and $\delta$, with larger $\omega$ or smaller $\delta$ increasing the number of operations. The flexibility of our implementation allows for easy adjustment of these parameters, enabling a balance between granularity of analysis and computational efficiency.

### v2

@@ III Related Work
TABLE I: COMPARISON OF RELATED WORKS ON INTRUSION DETECTION IN CANS.
[…]
@@ III-A Prior Work Closely Related to the Present Study
Islam et al. [Islam2020] proposed a four-stage IDS utilizing the chi-squared method to identify both strong and weak cyberattacks on the CAN bus. Their approach, which is among the first graph-based defense systems for CAN, showed misclassification rates of 5.26% for DoS attacks, 10% for fuzzy attacks, 4.76% for replay attacks, and zero misclassification for spoofing attacks. The dataset used for evaluation was obtained from the Hacking and Countermeasure Research Lab [Lee2018CANDataset]. This study emphasized the need for robust security mechanisms in modern vehicles and demonstrated superior accuracy compared to existing ID sequence-based methods.
@@
Jedh et al. [Jedh2021] proposed a message injection attack detection solution. Their approach leveraged MSGs and used graph-based analytics and anomaly detection to detect malicious message injections with high accuracy and low detection time. They validated their approach using a dataset collected from a moving vehicle under attack conditions. This study highlighted the importance of addressing the security of in-vehicle communication networks without relying on proprietary ECU information.
@@
Refat et al. [Refat2022] presented a lightweight IDS that translates CAN traffic into temporal graphs and applies neighborhood-based graph similarity techniques to detect message injection attacks. The system was evaluated using real vehicle data [Song2020InVehicleNetwork], achieving a detection accuracy of 96.01% for spoofing, fuzzy, and DoS attacks. This work contributed to vehicle security by providing a computationally efficient method without requiring changes to the CAN protocol.
@@
Zhang et al. [Zhang2023] proposed a CAN bus anomaly detection system using graph neural networks (GNNs) to detect message injection, suspension, and falsification attacks. Their approach involved creating directed attributed graphs from CAN message streams and training a two-stage classifier cascade. The evaluation on a Ford Transit 500 dataset [Othmane2022CANBusInjection] demonstrated the system’s efficiency in real-time detection and its ability to handle new anomalies through federated learning training.
@@
Xiao et al. [Xiao2023] proposed the CAN-GAT model based on graph attention networks for in-vehicle networks. By transforming CAN bus messages into graph structures, the model captured the correlation between traffic bytes, improving detection accuracy and efficiency. The model was evaluated using datasets from the Hacking and Countermeasure Research Lab [Lee2018CANDataset], demonstrating superior performance among compared GNNs.
@@
Meng et al. [Meng2023] developed GB-IDS, an IDS leveraging a novel graph structure and a variational autoencoder for training classifiers without negative samples. Their system, evaluated on the OTIDS dataset [Lee2017OTIDS], achieved high detection success rates for DoS, fuzzing, and impersonation attacks. This study addressed the limitations of traditional IDS, i.e., [Refat2022] by avoiding the need for protocol parsing and large training datasets.
Compared to prior works that focus on either graph-based or time series-based CAN IDS, our approach offers a unified framework that addresses their respective limitations. Graph-based methods, such as [Islam2020, Jedh2021], use MSGs to detect structural anomalies but lack granularity for subtle signal deviations. Time series-based techniques such as [moriano2024benchmarkingunsupervisedonlineids], [hanselmann2020canet] capture temporal anomalies but often miss intricate message sequence patterns and graph topology that are crucial for uncovering subtle masquerade attacks. These single-modality designs are insufficient for detecting stealthy masquerade attacks. Our framework combines graph embeddings and time series features, enabling detection across both structural and signal domains. Additionally, its flexibility—adjustable via window length and offset—supports fine-tuning in constrained environments. We evaluate this framework on the ROAD dataset [verma2022addressing], demonstrating improved performance over recent approaches [shahriar2022canshield, Zhang2023].
@@ III-B Other Prior Work Related to the Present Study
Hanselmann et al. [hanselmann2020canet] proposed CANet, a neural network architecture for detecting intrusions on the CAN bus. CANet, the first deep learning-based IDS capable of handling the high dimensionality of CAN bus data, was evaluated on real and synthetic CAN data. The method outperformed previous ML-based approaches with a high true negative rate, typically over 0.99, demonstrating robustness in detecting a large number of synthetic masquerade attacks [hanselmann2020canet].
Moriano et al. [Moriano:2022:Signal:Based:IDS] focused on detecting masquerade attacks on the CAN bus by analyzing time series extracted from raw CAN frames. Using hierarchical clustering, this study demonstrated that changes in cluster assignments could indicate anomalous behavior. The proposed forensic tool was tested on the ROAD dataset [verma2022addressing], showing significant differences in time series clustering similarity on benign and attack conditions.
Shahriar et al. [shahriar_cantropy] proposed CANtropy, a feature engineering-based lightweight CAN IDS. CANtropy explores a comprehensive set of features from both temporal and statistical domains utilizing PCA for anomaly detection. Evaluation on the SynCAN dataset [hanselmann2020canet] showed CANtropy’s effectiveness, with an average AUC-ROC score of 0.992, outperforming existing DL-based baselines.
Moriano et al. [moriano2024benchmarkingunsupervisedonlineids] conducted a benchmark study on four non-deep learning unsupervised online IDS for masquerade attacks in CANs. They controlled streaming data conditions in a sliding window setting and used realistic attacks from the ROAD dataset [verma2022addressing]. They noticed that there is no one-size-fits-all solution for detecting masquerade attacks in CAN in an online setting, i.e., there is no single algorithm that is optimal for detecting each attack type. Notably, a method that detects changes in the hierarchical structure of clusters of time series outperformed others in detecting various attack types at the expense of computational overhead.
@@ IV Proposed Data-Driven Methods
This section presents our approach for detecting masquerade attacks in CAN bus communication using graph ML. We begin by outlining the complete pipeline in Fig. 2, followed by a high-level algorithmic workflow in Algorithm 1. We then provide detailed procedures for MSG construction (Section IV-A), time series extraction (Section IV-B), node annotation (Section IV-C), graph embedding generation (Section IV-D), and supervised learning process (Section IV-E). Evaluation setup and metrics are described in Section IV-F.
@@
Threat model: In our framework, we assume the attacker has gained access to the CAN bus, either physically or remotely [checkoway2011comprehensive]. Masquerade attacks, as represented in the ROAD dataset [verma2022addressing], typically involve two compromised ECUs: a weak attacker that learns the frequency of messages sent by a targeted ECU and suppresses its transmissions, and a strong attacker that injects spoofed messages at the expected frequency using the same ID. The coordinated nature of this attack enables the adversary to mimic legitimate traffic patterns while altering the message content. In this study, our detection framework is evaluated under the assumption that the adversary can play both attacker roles, consistent with the ROAD dataset’s multi-stage masquerade attack model.
Algorithm 1 End-to-end masquerade attack detection
0: Raw CAN logs, $\omega_{t}$, $\delta_{t}$, DBC mappings, $node2vec_{params}$
0: Trained detection model
[…]
3:for each graph $G_{i}\in\mathcal{G}$ do
4:  $\text{AnnotateNodes}(G_{i},\mathcal{T})$ $\triangleright$ Alg. 3
5:  $E_{i}\leftarrow\text{node2vec}(G_{i},node2vec_{params})$ $\triangleright$ Sec. IV.D
6:  $F_{i}\leftarrow\text{Combine}(E_{i},G_{i}.\mu,G_{i}.\sigma)$ $\triangleright$ Feature fusion
7:end for
[…]
9:return $\mathcal{M}$
Note that to quantify the computational footprint of our end-to-end detection pipeline, we analyze its complexity.
Computational Complexity: The total cost is dominated by:
$\mathcal{T}_{total}$ $=\underbrace{O\left(\frac{n}{\delta_{t}}\cdot\omega_{t}\right)}_{\text{MSG building}}+\underbrace{O(n+k\cdot s_{sig})}_{\text{Time series processing}}$
$\quad+\underbrace{O(r\cdot l\cdot k)}_{\text{Graph embeddings}}+\underbrace{O(t\cdot d\cdot n_{train})}_{\text{ML Training}}$
Where:
  • $n$: Total CAN messages.
  • $k$: Unique CAN IDs.
  • $s_{\text{sig}}$: Maximum number of signals per CAN ID (not constant; varies across IDs, typically 2–10, with an upper limit of 16 in the ROAD dataset).
  • $n_{\text{train}}$: Number of training samples (approximately the number of sliding windows).
  • $\omega_{t}$, $\delta_{t}$: Window size and offset for time-based configurations (seconds); $\omega_{t}\in[2,15]$, $\delta_{t}\in[1,\omega_{t}]$.
  • $\omega_{s}$, $\delta_{s}$: Window size and offset for sample-based configurations (messages); $\omega_{s}\in\{50,100,\ldots,400\}$, $\delta_{s}\in\{50,100,\ldots,\omega_{s}\}$.
  • $r$: Number of node2vec walks per node (100).
  • $l$: Length of each node2vec walk (15).
  • $t$: Number of trees in the Random Forest classifier (100).
  • $d$: Maximum tree depth in the Random Forest classifier ($\approx\log_{2}n_{\text{train}}$)”.
Key Observations:
  • Training cost is negligible because:
    • It is done offline (not during vehicle operation)
    • Typically $n_{\text{train}}\ll 10^{4}$ windows
    • Random Forest scales linearly with $n_{\text{train}}$
  • Time series extraction is linear in $n$ (CAN-D decoding)
  • MSG construction dominates when $\omega_{t}\gg\delta_{t}$
  • Embedding generation (approximately 70% runtime) scales with $k^{2}$.
@@ IV-A CAN Data Partitioning and MSG Building
We use sliding windows of fixed size to partition and process the stream of CAN data [Qin2020]. The raw CAN data is first transformed into a structured dataframe format where each row represents a message with its timestamp, process ID, and data payload (see Fig. 3). This data is then partitioned into discrete time slices defined by specific window sizes and offsets, enabling a continuous analysis of CAN data.
@@
  • Window size ($\omega$, either $\omega_{t}$ for time or $\omega_{s}$ for samples): The length of each window determining the extent of data included in each graph representation.
  • Offset ($\delta$, either $\delta_{t}$ for time or $\delta_{s}$ for samples): The sliding step between consecutive windows, where $\delta$ represents the number of samples separating each window, ensuring no loss of data and the capture of communication patterns that may span multiple windows.
Parameter Rationale—Aligning with Attack Dynamics: The masquerade attacks in the ROAD dataset [verma2022addressing] are sustained, stealthy intrusions (e.g., correlated_signal, max_speedometer, max_engine_coolant_temp, reverse_light_off, reverse_light_on) lasting approximately 17–72 seconds. Our parameters were empirically optimized to capture attack duration while maintaining computational feasibility. For example, a 6-second window ensures sufficient temporal context to detect anomalies that unfold gradually, while a 3.2-second offset prevents missed detections near window boundaries through overlapping windows. This offset does not introduce a reaction delay but strategically balances attack coverage and computational feasibility—smaller offsets (e.g., 1s) increase processing load, whereas larger offsets (e.g., 6s) risk missing attacks. The selected window parameters ensure that our detection method aligns with real attack lifecycles.
Motivation for Sliding Window Detection: Masquerade attacks cannot be detected at the message level—they mimic legitimate traffic in short bursts but manifest as cumulative deviations over time. As demonstrated by Jedh et al. [Jedh2021], stealthy message injections, especially masquerade attacks, necessitate analysis over multiple consecutive frames rather than relying on single-message checks. Our approach aggregates e.g., 600 messages (at 100Hz) within a 6-second window, enabling structural (graph-based) and temporal (statistical) analysis to expose subtle attack patterns.
A MSG is constructed within each defined sliding window to model CAN activity following the procedure introduced by [Jedh2021] and [Islam2020]. Graphs are empirically recognized as effective tools to model relationships among relational data that are too complex for tabular representation or other simpler data structures [Bryan2023]. The process of building MSGs is crucial as it sets the stage for subsequent analyses, including node annotation and embedding generation. By modeling CAN data using MSG on sliding windows, we enhance our ability to detect and respond to potential masquerade attacks by finetuning the portion of analyzed data. We now detail each of the components of the MSG:
@@
  • Edges: Directed edges are established based on the sequential relationships of messages. That is, an edge is formed when one ID follows another, indicating the flow of communication [Islam2020]. This method effectively maps sequences of CAN messages enabling the detection of irregular patterns that may signify an attack. Edge weights are assigned to edges of MSGs patterns to quantify the frequency of observed communication sequences within a sliding window. Specifically, the weight of an edge from node $v_{i}$ to node $v_{j}$ is computed as the number of times the CAN message with ID $v_{j}$ directly follows the CAN message with ID $v_{i}$ within the same sliding window. Weights help to identify anomalous recurrent patterns that could indicate security threats.
Figure 3: Sliding window partitioning of CAN messages (left) and resulting MSG subgraph (right). Top five nodes by connectivity are highlighted; edges are weighted.
@@
0: Raw CAN data, $\omega_{t}$, $\delta_{t}$
@@
2: Initialize current_time=start_timecurrent\_time=start\_time of CAN data
3:while $current\_time+\omega_{t}\leq end\_time$ of CAN data do
4:  Define start_time=current_timestart\_time=current\_time and $end\_time=current\_time+\omega_{t}$
5:  Extract messages in [start_time,end_time][start\_time,end\_time]
@@
7:  for each message $m$ in [start_time,end_time][start\_time,end\_time] do
8:   Extract idid from $m$
9:   if idid not in $G_{sub}$ then
10:    Add idid as a node in $G_{sub}$
@@
12:   Determine next idid in sequence
13:   if edge from current idid to next idid exists then
@@
20:  Increment current_timecurrent\_time by $\delta_{t}$
@@
To ensure the correctness of Algorithm 2, we establish loop invariants for both the outer windowing loop and inner message processing loop. A loop invariant is a property that holds true before and after each iteration, ensuring algorithmic correctness. We validate these invariants through three steps:
@@
    Loop Invariants:
    • Outer Loop (Lines 3-21): At iteration $i$, the main graph $G$ contains all message sequences from windows $\{[t_{0},t_{0}+\omega_{t}],...,[t_{i-1},t_{i-1}+\omega_{t}]\}$ where $t_{j}=start\_time+j\cdot\delta_{t}$
    • Inner Loop (Lines 7-18): For message $m_{k}$ in window $t_{i},t_{i}+\omega_{t}]$, the subgraph $G_{sub}$:
      • Contains all nodes $\{id_{1},...,id_{k}\}$ from messages $\{m_{1},...,m_{k}\}$
      • Maintains edge weights equal to observed message sequence frequencies
@@
    Verification Steps:
    • Initialization:
      • Outer: $G=\emptyset$ before first window (trivially satisfies)
      • Inner: $G_{sub}=\emptyset$ before first message (trivially satisfies)
    • Maintenance:
      • Outer: Each iteration processes one complete window and merges $G_{sub}$ into $G$
      • Inner: Each message updates $G_{sub}$’s nodes/edges atomically
    • Termination:
      • Outer: Terminates when $current\_time+\omega_{t}>end\_time$ with all complete windows processed
      • Inner: Processes all messages in each window before continuing
  3. 3.
    Time Complexity:
    • Window processing:
      • $\lceil\frac{n}{\delta_{t}}\rceil$ windows processed
      • Each window requires $O(\omega_{t})$ operations (message extraction + graph updates)
    • Edge operations:
      • Average-case $O(1)$ for edge operations (networkx default implementation)
      • Each message generates exactly one edge update
      • Worst-case $O(k)$ per operation with hash collisions (kk=nodes)
    • Total complexity:
      • $O(\frac{n}{\delta_{t}}\cdot\omega_{t})$ for general case
      • $O(n)$ when $\omega_{t}=\delta_{t}$ (each message processed once)
Our experimental validation confirms this analysis:
  • Time-based windows: $\omega_{t}\in[2,15]$s, $\delta_{t}\in[1,15]$s (all combinations)
  • Sample-based windows: $\omega_{s}=\delta_{s}\in\{50k\mid k=1,\ldots,8\}$ messages
  • Linear scaling confirmed for $\omega/\delta\in[0.5,2.0]$